OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 11 –3x + 2y = 1

OpenStudy (anonymous):

Solution: We can add both of these equations together as follows: 3x + y = 11 -3x + 2y = 1 <-- imagine I put a huge addition sign to the very left of those equations.

OpenStudy (anonymous):

What would you get when you do (-3x) + (3x)? X cancels out! y + 2y = 3y 11 + 1 = 12 So now, what we are left with is this: 3y = 12 Solving for y, y = 4.